Evaluate as instructed. 1) Use f(x) = 3x + 4 and g(x) = x - x2 to evaluate (f + g)(-2). A) -4 C-8 B) 4 D) 12
Answer #1

0.0 Fox)- 3x+4, 966) = *-* (feg) 6) = f(x) eg(x) = 3x+4+x-x? fega) = -x244844 :: Et9) (-2)=-(-2)? 4 (-2) +4 --4 -8 +4 : feg)(